### v1

@@ IV.3 Theoretical studies
Nuclear reactions of astrophysical interest in general, and the 2H(,)3He in particular, are of great importance in nuclear theory because the available experimental data can be used to test the adopted theoretical framework. The 2H(,)3He reaction has the great advantage of involving only nuclear systems, and can be addressed with a microscopic ab initio study. This means that the nuclear systems involved in the process are viewed as made up of nucleons, interacting among themselves and with the external electromagnetic probes. Within such an approach, the following ingredients are essential for the calculation: (i) realistic models for the nuclear interactions and currents, possibly rooted in Quantum Chromodynamics (QCD); (ii) a numerical technique able to solve the -body bound and scattering state problem, including the Coulomb interaction without approximation. Such a technique is usually referred to as an ab initio method. The agreement (or disagreement) between ab initio theoretical predictions and experimental data represents a validation (or indicates the necessity of improvement) mostly for ingredient (i), i.e. the models of the nuclear interactions and currents, within the accuracy of the ab initio method, i.e., ingredient (ii). This is why few-nucleon reactions can be used as an “ideal” laboratory, where the ab initio framework can be stringently tested.
@@ IV.4 Phenomenological and Bayesian Analyses
In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the factor can be modelled by a low-order polynomial in energy [1967ARA&A...5..525F]. Moreover, the recent results from LUNA [2020Natur.587..210M], combined with previous measurements, place stringent constraints on the cross section such that a rigorous analysis can capture both the average value and uncertainty over the range of interest to Solar fusion and Big Bang Nucleosynthesis [2020Natur.587..210M, 2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y].
@@
Another change since SF-II has been the advent of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [2016ApJ...831..107I]. To describe this approach, we closely follow the analysis of the 2H(,)3He reaction in 2021ApJ...923...49M, and references therein.
Consider an optimization that utilizes Gaussian distributed priors in a constrained Bayesian analysis, such that the augmented (for uncorrelated data) is given by
The first term in the above equation is the standard and the second term derives from the Bayesian constraint on the parameters with prior mean () and width (), which are chosen with some prior knowledge or can be optimized as described below. In the first term, the double sum runs over the datasets () and individual results from each dataset (), with the mean value and stochastic uncertainty of each data point given by and , respectively.
The theoretical model describing the factor data, , is a function of the energy () and a set of parameters () that must be determined. The quoted systematic uncertainties are parameterized by the normalization factors , with a prior of unit normalization and a width characterized by the quoted systematic uncertainty. Within a Bayesian framework, it is straightforward to utilize distribution functions for that are not Gaussian, such as a log-normal or other distributions. The parameters represent a normalization of the model function for a given dataset, which, viewed from a non-Bayesian perspective, can also be interpreted as normalization factors that must be applied to the data to match the “true” underlying distribution.
Finally, are unknown extrinsic uncertainties [2019PhRvC..99a4619D]. If one assumes that a smooth function of energy can accurately describe the data, then the data in a given experimental set may scatter about this presumed “true” value by more than is reflected by the quoted statistical uncertainties. It was suggested in 2019PhRvC..99a4619D that this extra scatter might be explained by some additional source of statistical uncertainty unbeknownst to the experimenter, or by some unknown systematic uncertainty that is different for each data point in the same data set, as opposed to a correlated systematic that affects all data points similarly. In either case, this extrinsic uncertainty can be accommodated in a Bayesian analysis framework by adding the additional uncertainty as a normal-distributed source of noise with a width that is constrained by the data. In 2019PhRvC..99a4619D and 2021ApJ...923...49M, was added as an absolute uncertainty, independent of the energy. For some data, 2022PhRvC.105a4625O suggested adopting instead a relative uncertainty, such that the scale of the extrinsic fluctuations is proportional to the mean value of . This strategy of adding extra extrinsic uncertainty to the data sets can be viewed as an alternative to that of inflating the quoted statistical uncertainties by , which is often used for seemingly incompatible data sets [2022PTEP.2022h3C01W], where is the per degree of freedom. An advantage of the extrinsic uncertainty method is that it uses the observed scatter within a given data set as a measure of the possible size of unreported uncertainties, as opposed to uniformly increasing the uncertainty in all data sets by the same relative amount.
@@
Such a Bayesian analysis of the 2H(,)3He reaction data was performed by 2021ApJ...923...49M, using both a third-order polynomial and the ab initio prediction from 2005PhRvC..72a4001M. As noted in the literature, the energy dependence of 2005PhRvC..72a4001M is more reliable than the absolute normalization. Furthermore, the updated prediction in 2016PhRvL.116j2501M, as noted above, is larger than that of 2005PhRvC..72a4001M. Therefore, in 2021ApJ...923...49M, the ab initio prediction is modeled as
where is the prediction from 2005PhRvC..72a4001M (or alternatively from 2016PhRvL.116j2501M), and and are an unknown scale factor and offset to be determined in the analysis. While was not determined with theoretical uncertainty, model uncertainty is introduced through the parameters and . The resulting mean values and uncertainties of determined from Eq. (11) were found to be comparable. For a given model, after optimizing the posterior parameter distributions, the Baye factor (BF) is proportional to the probability of the model given the data [2017bmad.book.....H]. Therefore, for a fixed data set, the BF can be used as a relative probability of each model, thus enabling a weighted model-averaging procedure. If we assume a uniform likelihood for each model, the expectation value and variance of a quantity is given by
where denotes the expectation of given the model and denotes the probability of the model given the data (), which is given by
with the Baye factor of model .
Similarly, for a given model, in the absence of prior information on the size of an unknown parameter, the optimal width of its prior can be estimated by finding the value of that maximizes the BF, which typically provides a reasonable approximation to marginalizing over the prior width.
| keV | 10 | 20 | 40 | 80 | 91 | 100 | 120 |
@@ IV.5 Summary and Recommendations
With such a Bayes model averaging, we can compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by 2005PhRvC..72a4001M, 2016PhRvL.116j2501M, including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq. (11). Results are listed in Table 3. The first column lists the model (i.e., polynomial or ab initio), the second column the natural logarithm of the Gaussian approximation to the BF, and the third column the corresponding weight in the model averaging. From this model averaging, the resulting prediction for the 2H(,)3He factor at a few representative energies is provided in Table 2, where the first uncertainty arises from the first term in Eq. (IV.4) and the second uncertainty is from the second term, which we denote as model selection uncertainty. The resulting Bayes model average prediction of the factor over the entire kinematic range considered is depicted in Fig. 5, with the gray band representing the 68% coverage probability 555The analysis in Table 3 and Fig. 5 can be reproduced with the code at the git repository https://github.com/nrp-g/leaner.. For this particular reaction, it is interesting to note that the Bayesian analysis strongly favors the phenomenological models of Eq. (11) over the polynomial approximations. One reason for this might be that the scaled ab initio models have only one or two free parameters, and the energy dependence of given in 2005PhRvC..72a4001M, 2016PhRvL.116j2501M is sufficient to accurately describe the various data sets. On the other hand, polynomial approximations are disfavored as they require more parameters to capture the energy dependence. Also of note, in the simplest model, using 2016PhRvL.116j2501M, the scale factor is given by
@@
indicating that the prediction in 2016PhRvL.116j2501M overestimates the 2H(,)3He data by 7.9%, consistent with expectations. In comparison, the third order polynomial fit predicts values of that are higher at and lower at keV as compared to those in Table 2.
V The reaction ()
@@
The 3He(3He,)4He reaction (astrophysical factor ) terminates the pp-I chain. The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in SF-I increasing was discussed as a potential solution to the Solar Neutrino Problem. Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [1999PhRvL..82.5205B] and a complementary experiment at somewhat higher energies [2004PhRvC..69a5802K] ruled out such an increase in , as summarized in SF-II.
@@ V.1 Shape of the particle spectrum
Since no new absolute measurements of have been reported since SF-II, we consider the same four experiments [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K]. However, there is new information on the energy spectrum of the emitted protons. This spectrum has recently been measured using Inertial Confinement Fusion (ICF) plasmas for a Gamow peak energy of 165 keV [2017PhRvL.119v2701Z]. The results show significant structure, indicating the presence of a sequential reaction mechanism passing through the unbound ground state of . This spectrum is important for the measurements [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K] since they determined cross sections by detecting only the protons above an energy threshold.
@@ V.2 Recommendation
Based on these investigations, our recommended fit is the one from SF-II, but with an additional 4% systematic uncertainty. This gives
Only the larger constant term in the uncertainty differs from the previous evaluation.
@@
VI The reaction ()
The 3He()7Be reaction, or , proceeds via direct capture to the ground and 429 keV first excited states of 7Be. It has been studied experimentally using three main methods: First, by detecting at least two of the three -rays from the reaction and taking the angular correlation with the alpha beam direction into account (prompt- method). Second, by detecting the induced 7Be [2002NuPhA.708....3T] radioactivity (activation method). Third, by counting the 7Be recoils (recoil method).
@@ VI.1 Previous recommendation in SF II
The 3He()7Be factor recommended in SF II was developed in multiple steps. First, a model was selected for the shape of , based on existing nucleon-level calculations [1986NuPhA.460..559K, 2001PhRvC..63e4002N]. A rescaling of the curve was fitted to data, but only at 1.002 MeV center-of-mass to minimize the role of short-distance physics in the models.
In SF I, a possible systematic discrepancy between data from the two methods previously used was discussed. By the time of SF II, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] and Seattle [2007PhRvC..76e5801B] groups. These studies did not find any discrepancy between activation and prompt- data in direct comparisons. SF II opted to limit the fitting to the data by the activation and recoil methods. The prompt- data were left out because of their somewhat larger common mode errors, concerns about how well the -ray angular distribution was known, and avoidance of the correlated errors between activation and prompt data from the same experiment.
In SF II, only data published after 1998 were included. This was partly because the newer data are in general better documented than the older data, and partly because the newer experiments were perceived as addressing systematics more quantitatively. Data from each of the retained experiments were used to determine a separate rescaling of the theory curve, and the rescalings were averaged to find a recommended and recommended error. Theory error was estimated by dispersion of the extrapolated among several models, including versions of the main fitting model that were modified to cover the experimentally-allowed range of scattering lengths.
The SF II recommended value was , based on the data by the Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G], Seattle [2007PhRvC..76e5801B], and ERNA [2009PhRvL.102w2502D] groups.
@@ VI.2 Theory progress on
Significant theoretical work on this reaction has occurred since SF II, but the basic understanding of its mechanism remains unchanged from the 1960s [1963PhRv..131.2582T]. It is dominated by external direct capture into the two bound states, and most of the dipole strength at low energy arises beyond the range of nuclear interaction. The external capture part of the cross section is determined by the ANCs of the bound states and by scattering phase shifts; near threshold, most of the strength lies at –20 fm [2011PhRvL.106d2502N]. In models with explicit wave functions the shorter-range strength largely cancels out due to effects of nucleon-exchange antisymmetry. All models feature a shallow minimum of near 1.25 MeV, where capture from -waves becomes comparable to that from -waves.
The first fully ab initio calculation [2011PhRvL.106d2502N] appeared just after the SF II analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [1995PhRvC..51...38W]. It agrees well with both the scale and the energy dependence of the modern factor data, and it is very close to the energy dependence assumed in SF II. Neff’s elastic-scattering phase shifts also agree well with experiment. Notably, Neff found that the dipole strength distribution departs significantly from pure external capture at 3He-4He separations as large as 9 fm, compared with 3-4 fm in potential models.
@@
An important advance since SF II has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [2002NuPhA.712...37B, 2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. EFTs are valid for systems with a natural separation between the momentum scales probed in low-energy experiment (e.g., corresponding to a binding energy or a projectile energy) and the much larger momentum scale where the low-energy degrees of freedom are no longer valid: the word “halo” here refers to nuclei with only shallow bound states. In halo EFT the only degrees of freedom are the initial- and final-state nuclei treated as point particles, plus photons. For , MeV corresponds to the momentum needed to separate a proton from 3He or to excite 4He; the thresholds to separate the two 7Be bound states into 3He and 4He correspond to momenta MeV and 60.9 MeV.
@@
Two groups have studied in halo EFT [2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. The sizes of the Coulomb interaction and the large -wave scattering length make the correct power counting tricky to establish for this system. The work of 2018EPJA...54...89H and 2020EPJA...56..166P includes a careful examination of possible power-counting schemes and strict adherence to a power counting once established, up to terms of NLO or NNLO, respectively, in two different power-countings. These authors fitted EFTs to factor data, and they examined effects of including or excluding elastic scattering constraints in their fits. When they included scattering phase shifts from 1972NuPhA.195..241B in their fits, they concluded that the large-scattering-length power counting of the NLO theory was favored. This result promotes two-body currents (i.e., contributions not equivalent to external direct capture) to leading order in the theory. The -wave contribution that becomes important above 1 MeV first appears at NLO. In addition to data-fitting errors, these theories have errors estimated to be 10% from EFT truncation.
@@
The work of 2020JPhG...47e4002Z followed a different approach to power counting up through NLO. At MeV, their derived expression for is essentially the same as that from the 2020EPJA...56..166P NLO theory. 2020JPhG...47e4002Z also developed some ad hoc (i.e., not systematically developed) higher-order EFT terms, referred to as partial-N4LO, to test for their impact on the fitting. The additional terms proved not to be required by the data and did not improve the fit; this result was taken to indicate that corrections from omitted terms are not large compared with experimental errors below 2 MeV. Scattering data were not considered in this work apart from very broad priors on scattering length and effective range; the correlated errors in the factor data were taken into account.
@@
Like halo EFT, the phenomenological R-matrix approach avoids a model of nuclear interactions and uses a systematic parameterization to fit data [2010RPPh...73c6301D]. At the time of SF II, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in 2004ADNDT..88..203D that focused on BBN energies. After SF-II an R-matrix analysis using the AZURE2 code [2010PhRvC..81d5805A, azure2-manual] was carried out on both elastic scattering and factor data in conjunction with the Notre Dame experiment [2013PhRvC..87f5804K], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in 2014PhRvC..90c5804D. In the latter analysis both the fitted value and error bars were heavily influenced by the numerous scattering data of 1964NucPh..50..629B. Very recently, the BRICK software package has been constructed to carry out Bayesian parameter estimation for AZURE2 and applied to both capture and elastic scattering at all energies in the 7Be system by 2022FrP....10.8476O. In this work it was shown that markedly different values result from inclusion or not of the older scattering data [1964NucPh..50..629B] alongside the very recent SONIK scattering data of 2024PhRvC.109a5802P.
@@
A small amount of additional theoretical work in more traditional frameworks has appeared since SF-II. This includes fits of potential models to factor data [2018PhRvC..97c5802T, 2021NuPhA100622108T] and some calculations of the RGM type [2017PhRvC..96f4605S, 2019PhRvC..99e4618S].

### v2

@@ IV.3 Theoretical studies
Nuclear reactions of astrophysical interest in general, and the 2H(,)3He in particular, are of great importance in nuclear theory because the available experimental data can be used to test the adopted theoretical framework. The 2H(,)3He reaction has the great advantage of involving only nuclear systems, and can be addressed with a microscopic ab initio study. This means that the nuclear systems involved in the process are viewed as made up of nucleons, interacting among themselves and with the external electromagnetic probes. Within such an approach, the following ingredients are essential for the calculation: (i) realistic models for the nuclear interactions and currents, possibly rooted in Quantum Chromodynamics (QCD); (ii) a numerical technique able to solve the -body bound and scattering state problem, including the Coulomb interaction without approximation. Such a technique is usually referred to as an ab initio method. The agreement (or disagreement) between ab initio theoretical predictions and experimental data represents a validation (or indicates the necessity of improvement) mostly for ingredient (i), i.e. are the models of the nuclear interactions and currents accurate within the precision specified by the ab initio method? This is why few-nucleon reactions can be used as an “ideal” laboratory, where the ab initio framework can be stringently tested in systems under more theoretical control.
@@ IV.4 Phenomenological and Bayesian Analyses
In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the S-factor can be modelled by a low-order polynomial in energy [1967ARA&A...5..525F]. Moreover, the recent results from LUNA [2020Natur.587..210M], combined with previous measurements, place stringent constraints on the cross section, reducing the uncertainty over the range of interest to Solar fusion and BBN [2002NuPhA.706..203C, 2020Natur.587..210M, 2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y].
@@
A significant change since 2011RvMP...83..195A has been the wide adoption of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [2016ApJ...831..107I]. This approach has been followed in the analysis presented here, which closely follow work on the 2H(,)3He reaction reported in 2021ApJ...923...49M, and references therein. The mathematical details of the Bayesian method will be found in Appendix LABEL:AppendixWG4: Here we focus on the application to 2H(,)3He by the authors cited above.
Both the ab initio prediction from 2005PhRvC..72a4001M as well as a third-order polynomial in the energy were used to constrain the data, where we define an -order polynomial as
@@
As noted in the literature, the energy dependence of 2005PhRvC..72a4001M is more reliable than the absolute normalization. Furthermore, the updated prediction in 2016PhRvL.116j2501M, as noted above, is larger than that of 2005PhRvC..72a4001M. Therefore, 2021ApJ...923...49M modeled the prediction of 2005PhRvC..72a4001M, denoted , as
where and are an unknown scale factor and offset to be determined in the analysis. While was not determined with theoretical uncertainty, model uncertainty is introduced through the parameters and . The resulting mean values and uncertainties of determined from Eq. (11) were found to be comparable to those with the third-order polynomial.
@@ IV.5 Summary and Recommendations
In this review, we take the further step of applying Bayes Model Averaging to a number of reasonable models that describe the data, which systematically captures additional theoretical “model selection uncertainty”. Detailes of the Bayesian analysis and model averaging procdure are presented in Appendix LABEL:AppendixWG4. The models we explore include Eq. (11) with both and , as done in 2021ApJ...923...49M, as well as setting , and then we consider polynomials in energy of order .777The implementation of the Bayesian analysis presented here is different than that of 2021ApJ...923...49M but it was verified to produce the same results when the same implementation of the systematic uncertainties was made.
@@
With such a Bayes model averaging, we can quantitatively compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by 2005PhRvC..72a4001M, 2016PhRvL.116j2501M, including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq. (11). Results are listed in Table 3. The first column lists the model (i.e., polynomial or ab initio), the second column the natural logarithm of the Gaussian approximation to the BF, and the third column the corresponding weight in the model averaging. From this model averaging, the resulting prediction for the 2H(,)3He S-factor at a few representative energies is provided in Table 4, where the first uncertainty arises from the first term in Eq. (LABEL:eq:BMA_Var) and the second uncertainty is from the second term, which we denote as model selection uncertainty. The resulting Bayes model average prediction of the S-factor over the entire kinematic range considered is depicted in Fig. 5, with the gray band representing the 68% coverage probability. For this particular reaction, it is interesting to note that the Bayesian analysis strongly favors the phenomenological models of Eq. (11) over the polynomial approximations. One reason for this might be that the scaled ab initio models have only one or two free parameters, and the energy dependence of given in 2005PhRvC..72a4001M, 2016PhRvL.116j2501M is sufficient to accurately describe the various data sets. On the other hand, polynomial approximations are disfavored as they require more parameters to capture the energy dependence. Also of note, in the model of Eq. (11) with and using 2016PhRvL.116j2501M for , the scale factor is given by
@@
indicating that the prediction in 2016PhRvL.116j2501M overestimates the 2H(,)3He data by 7.9%, consistent with expectations noted above. In comparison, the third order polynomial fit predicts values of that are higher at and lower at keV as compared to those in Table 4.
The analysis in Table 3 and Fig. 5 can be reproduced with the code linked in Table 4. The code also can provide a prediction of the mean value and uncertainty of at any energy over the same kinematic range.
| keV | 10 | 20 | 40 | 80 | 91 | 100 | 120 |
The reaction )
@@
The 3He(3He,)4He reaction terminates the pp-I chain. The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in 1998RvMP...70.1265A increasing was discussed as a potential solution to the Solar Neutrino Problem. Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [1999PhRvL..82.5205B] and a complementary experiment at somewhat higher energies [2004PhRvC..69a5802K] ruled out such an increase in , as summarized in 2011RvMP...83..195A.
@@ V.1 Shape of the particle spectrum
Since no new absolute measurements of have been reported since 2011RvMP...83..195A, we consider the same four experiments [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K]. However, there is new information on the energy spectrum of the emitted protons. This spectrum has recently been measured using Inertial Confinement Fusion (ICF) plasmas for a Gamow peak energy of 165 keV [2017PhRvL.119v2701Z]. The results show significant structure, indicating the presence of a sequential reaction mechanism passing through the unbound ground state of . This spectrum is important for all of the measurements since they determined cross sections by detecting only the protons above an energy threshold.
@@ V.2 Recommendation
Based on these investigations, our recommended fit is the one from 2011RvMP...83..195A, but with an additional 4% systematic uncertainty typical of the 3% to 6% corrections to the extrapolated proton spectra estimated above to take the uncertain spectral shape into account. This gives
Only the larger constant term in the uncertainty differs from the previous evaluation, and it is due to the spectral shape discussed above.
@@
The reaction ()
The 3He()7Be reaction proceeds via non-resonant capture to the ground and 429 keV first excited states of 7Be. It has been studied experimentally using three main methods, with references given in subsections VI.1 and VI.3: First, by detecting at least two of the three -rays from the reaction and taking the angular correlation with the alpha beam direction into account (prompt- method). Second, by detecting the induced 7Be [2002NuPhA.708....3T] radioactivity (activation method). Third, by counting the 7Be recoils (recoil method).
@@ VI.1 Previous recommendation in SF II
The 3He()7Be S-factor recommended in 2011RvMP...83..195A was developed in multiple steps. First, a model was selected for the shape of , based on existing nucleon-level calculations [1986NuPhA.460..559K, 2001PhRvC..63e4002N]. A rescaling of the curve was fitted to data, but only at 1.002 MeV center-of-mass to minimize the role of short-distance physics in the models.
In 1998RvMP...70.1265A, a possible systematic discrepancy between data from the two methods previously used was discussed. By the time of 2011RvMP...83..195A, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] and Seattle [2007PhRvC..76e5801B] groups. These studies did not find any discrepancy between activation and prompt- data in direct comparisons. 2011RvMP...83..195A opted to limit the fitting to the data by the activation and recoil methods. The prompt- data were left out because of their somewhat larger common mode errors, concerns about how well the -ray angular distribution was known, and avoidance of the correlated errors between activation and prompt data from the same experiment.
In 2011RvMP...83..195A, only data published after 1998 were included. The 2011RvMP...83..195A recommended value was , based on the data by the Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G], Seattle [2007PhRvC..76e5801B], and ERNA [2009PhRvL.102w2502D] groups.
@@ VI.2 Theory progress on
Significant theoretical work on this reaction has occurred since 2011RvMP...83..195A, but the basic understanding of its mechanism remains unchanged from the 1960s [1963PhRv..131.2582T]. It is dominated by external non-resonant capture into the two bound states, and most of the dipole strength at low energy arises beyond the range of nuclear interaction. The external capture part of the cross section is determined by the ANCs of the bound states and by scattering phase shifts; near threshold, most of the strength lies at –20 fm [2011PhRvL.106d2502N]. In models with explicit wave functions the shorter-range strength largely cancels out due to effects of nucleon-exchange antisymmetry. All models feature a shallow minimum of near 1.25 MeV, where capture from -waves becomes comparable to that from -waves.
The first fully ab initio calculation [2011PhRvL.106d2502N] appeared just after the 2011RvMP...83..195A analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [1995PhRvC..51...38W]. It agrees well with both the scale and the energy dependence of the modern S-factor data, and it is very close to the energy dependence assumed in 2011RvMP...83..195A. Neff’s elastic-scattering phase shifts also agree well with experiment. Notably, Neff found that the dipole strength distribution departs significantly from pure external capture at 3He-4He separations as large as 9 fm, compared with 3-4 fm in potential models.
@@
An important advance since 2011RvMP...83..195A has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [2002NuPhA.712...37B, 2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. EFTs are valid for systems with a natural separation between the momentum scales probed in low-energy experiment (e.g., corresponding to a binding energy or a projectile energy) and the much larger momentum scale where the low-energy degrees of freedom are no longer valid: the word “halo” here refers to nuclei with only shallow bound states. In halo EFT the only degrees of freedom are the initial- and final-state nuclei treated as point particles, plus photons. For , MeV corresponds to the momentum needed to separate a proton from 3He or to excite 4He; the thresholds to separate the two 7Be bound states into 3He and 4He correspond to momenta MeV and 60.9 MeV.
@@
Two groups have studied in halo EFT [2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. The sizes of the Coulomb interaction and the large -wave scattering length make the correct power counting tricky to establish for this system. The work of 2018EPJA...54...89H and 2020EPJA...56..166P includes a careful examination of possible power-counting schemes and strict adherence to a power counting once established, up to terms of NLO or NNLO, respectively, in two different power-countings. These authors fitted EFTs to S-factor data, and they examined effects of including or excluding elastic scattering constraints in their fits. When they included scattering phase shifts from 1972NuPhA.195..241B in their fits, they concluded that the large-scattering-length power counting of the NLO theory was favored. This result promotes two-body currents (i.e., contributions not equivalent to external non-resonant capture) to leading order in the theory. The -wave contribution that becomes important above 1 MeV first appears at NLO. In addition to data-fitting errors, these theories have errors estimated to be 10% from EFT truncation.
@@
The work of 2020JPhG...47e4002Z followed a different approach to power counting up through NLO. At MeV, their derived expression for is essentially the same as that from the 2020EPJA...56..166P NLO theory. 2020JPhG...47e4002Z also developed some ad hoc (i.e., not systematically developed) higher-order EFT terms, referred to as partial-N4LO, to test for their impact on the fitting. The additional terms proved not to be required by the data and did not improve the fit; this result was taken to indicate that corrections from omitted terms are not large compared with experimental errors below 2 MeV. Scattering data were not considered in this work apart from very broad priors on scattering length and effective range; the correlated errors in the S-factor data were taken into account.
@@
Like halo EFT, the phenomenological R-matrix approach avoids a model of nuclear interactions and uses a systematic parameterization to fit data [2010RPPh...73c6301D]. At the time of 2011RvMP...83..195A, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in 2004ADNDT..88..203D that focused on BBN energies. After 2011RvMP...83..195A an R-matrix analysis using the AZURE2 code [2010PhRvC..81d5805A, azure2-manual] was carried out on both elastic scattering and S-factor data in conjunction with the Notre Dame experiment [2013PhRvC..87f5804K], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in 2014PhRvC..90c5804D. In the latter analysis both the fitted value and error bars were heavily influenced by the numerous scattering data of 1964NucPh..50..629B. Very recently, the BRICK software package has been constructed to carry out Bayesian parameter estimation for AZURE2 and applied to both capture and elastic scattering at all energies in the 7Be system by 2022FrP....10.8476O. In this work it was shown that markedly different values result from inclusion or not of the older scattering data [1964NucPh..50..629B] alongside the very recent SONIK scattering data of 2024PhRvC.109a5802P.
@@
A small amount of additional theoretical work in more traditional frameworks has appeared since 2011RvMP...83..195A. This includes fits of potential models to S-factor data [2018PhRvC..97c5802T, 2021NuPhA100622108T] and some calculations of the RGM type [2017PhRvC..96f4605S, 2019PhRvC..99e4618S].